# Spurious Correlation-Aware Embedding Regularization for Worst-Group Robustness

### v1

@@ B.1 Experimental Setting
This section provides comprehensive details of our experimental configuration for evaluating the SCER framework. We describe the datasets used for evaluation, model architectures employed across different domains, baseline methods for comparison, hyperparameter settings, and evaluation metrics. These detailed specifications ensure reproducibility and provide a complete context for the results presented in the main paper.
@@ B.1.1 Datasets
We conduct experiments on six benchmark datasets: four image datasets (Waterbirds, CelebA, MetaShift, ColorMNIST) and two text datasets (CivilComments, MultiNLI). These datasets are established benchmarks for evaluating model robustness against spurious correlations. Dataset characteristics and statistics are summarized in Table 8.
Table 8: Statistics of datasets. For ColorMNIST, $\rho$ denotes the spurious correlation level between label $y$ and domain attribute (color) in the training set. A higher $\rho$ indicates stronger spurious correlation, making color a more predictive but misleading cue for classification. "One Subpopulation Omitted" represents an extreme case where one group combination is completely absent from training data. Groups are defined as combinations of class labels and attribute values, and Max/Min group columns show the sample sizes of the largest and smallest groups, respectively. The disparity between these values indicates the degree of group imbalance, which tests model robustness under uneven data distribution.
@@
Waterbirds (sagawadistributionally) is a binary classification dataset constructed By superimposing bird images from the CUB dataset (wah2011caltech) onto backgrounds from the Places dataset (zhou2017places). The task involves classifying images as landbirds or waterbirds, where background type creates spurious correlations between land backgrounds and landbirds, and water backgrounds and waterbirds. We use the standard data splits from idrissi2022simple.
@@
CelebA (liu2015faceattributes) contains approximately 200,000 celebrity images. We formulate the problem as a binary classification, classifying hair as either blond or not blond, with gender serving as the spurious attribute. We follow the standard splits from idrissi2022simple.
@@
From the MetaShift benchmark (liang2022metashift), we use the Cats vs. Dogs binary classification task, where background type serves as the spurious attribute through correlations between indoor backgrounds with cats and outdoor backgrounds with dogs. We adopt the unmixed version provided by the dataset authors.
@@
ColorMNIST (arjovsky2019invariant) is a synthetic variant of MNIST where each digit is assigned a color. The classification task groups digits into two categories: zero and four versus five and nine, with color serving as the spurious attribute through red and green assignments. Training data exhibits strong color-class correlations, validation data is balanced at 50:50, and test data reverses the correlation pattern to 10:90. Following arjovsky2019invariant, we inject 25% label noise and evaluate under varying spurious correlation strengths of 80:20, 90:10, and 95:5 ratios to assess robustness.
To test the limits of group imbalance, we conducted additional experiments on ColorMNIST with severe distributional skew where one subpopulation is completely omitted from training. Specifically, we use 15,249 samples for Label 0 with Color 0, 3,000 samples for Label 1 with Color 0, and 11,751 samples for Label 1 with Color 1, while Label 0 with Color 1 contains 0 samples. This extreme setting evaluates model performance when certain subpopulations are entirely missing during training, representing a challenging scenario for robust classification methods.
To test the generalizability of SCER, we conducted additional experiments on ColorMNIST using the environment inference methodology from creager2021environment. Following the prior protocol with two environments in train, we partition the dataset into 50,000 training samples and 10,000 test samples. The training set is further divided into two environments, each containing 25,000 samples, representing different spurious correlation patterns. We maintain consistency with our primary experimental setup by using identical model architectures, training procedures, and evaluation metrics across all experiments.
@@
CivilComments (borkan2019nuanced) provides a binary text classification benchmark designed to identify toxic language in online comments. Mentions of demographic groups serve as spurious attributes. We use the standard WILDS splits.
@@
The MultiNLI dataset (williams2017broad) addresses natural language inference by requiring models to classify the relationship between a premise and a hypothesis as entailment, contradiction, or neutrality. The presence of negation words serves as a confounding factor because they exhibit a strong correlation with contradiction labels. We adopt the data splits established in idrissi2022simple.
B.1.2 Baselines
For fair comparison, we followed the training protocol of gulrajani2020search, performing random search over each method’s joint hyperparameter space. Hyperparameter configurations follow prior work (yang2023change) for all baseline methods. We selected the hyperparameters that achieved the highest Worst Accuracy for each algorithm, then conducted three independent runs with different random seeds (0, 1, 2) to account for variance and ensure robustness of reported results. Final average performance and standard deviation are reported across these runs. Complete hyperparameter details are provided in Table 9 summarizes the hyperparameters used for each baseline. In addition to standard optimization parameters such as batch size, learning rate, and dropout, we include key weighting parameters $\eta$, $\lambda_{spur}$, and $\lambda_{core}$, which are used to suppress spurious features and enhance core features. These values are chosen to encourage the model to learn generalizable representations under distribution shifts. To comprehensively evaluate the robustness of our method, we compare its performance against a wide range of baselines, including standard ERM, subpopulation robustness methods, domain-invariant approaches, data augmentation techniques, and class imbalance methods.
@@ B.1.2 Baselines
GroupDRO (sagawadistributionally) is a representative approach to distributional robustness that improves worst-group performance by assigning greater importance to groups with higher losses. CVaRDRO (duchi2021learning) extends this approach to the individual sample level, enabling more fine-grained adjustment by assigning importance weights to individual samples. LfF (nam2020learning) employs a co-training approach that simultaneously trains a biased model and a debiased model, guiding the model to learn by distinguishing between biased and unbiased features. JTT (pmlr-v139-liu21f) is a two-stage strategy that identifies hard samples using information obtained from initial standard empirical risk minimization (ERM) training, then improves model robustness by increasing the weights of these samples in the second training stage. LISA (yao2022improving) utilizes data augmentation techniques that mix data within and across attributes, encouraging the model to learn invariant features of attributes. DFR (kirichenko2022last) is an efficient method that improves robustness by retraining only the last layer using a balanced validation set, rather than the entire model. PDE (deng2023robust) employs a strategy of progressively expanding training data, guiding the model to first focus on core features and subsequently learn spurious features. Finally, ElRep (wen2025elastic) introduces both Nuclear-norm and Frobenius-norm regularization to representation vectors, enabling robust group generalization representation learning against spurious correlations by simultaneously emphasizing important features and maintaining diversity.
@@
IRM (arjovsky2019invariant) is a method for finding invariant predictors that work commonly across multiple domains. Specifically, it trains feature extractors such that a fixed predictor becomes optimal in each domain, enabling the model to effectively capture important characteristics regardless of domain changes. This approach ensures robust performance even when domains change. MMD (li2018domain) is a method that numerically measures the difference in feature distributions between different domains, training to reduce this difference to mitigate distribution mismatches across domains. In other words, it is an approach that aims to improve generalization performance by making feature representations across domains as similar as possible.
@@
Mixup (zhang2018mixup) is a technique that creates new synthetic training data by linearly interpolating two random samples and their labels. This smooths decision boundaries and serves as regularization to improve generalization performance and prevent overfitting. It is a simple yet effective data augmentation method that artificially expands training data diversity by learning intermediate points between samples in the data space.
@@
ReSample (japkowicz2000class) mitigates class imbalance in training data through resampling techniques. ReWeight (japkowicz2000class) assigns weights to the loss function proportional to class frequency. SQRTWeight is another loss adjustment method that modifies loss weights using the inverse square root of class frequency, appropriately increasing the influence of minority classes in imbalanced data. Focal Loss (lin2017focal) focuses on difficult samples by reducing the contribution of easy-to-learn samples. CB Loss (cui2019class) calculates sample weights based on effective class size. LDAM (cao2019learning) assigns different margins for each class to balance decision boundaries. Balanced Softmax (ren2020balanced) modifies the softmax function itself to accommodate class imbalance.
We compared our model against these numerous baselines and demonstrated the superiority of our approach, confirming the necessity of embedding regularization.
@@ C.1 Quantitative Analysis
We conduct a comprehensive quantitative analysis to evaluate the computational complexity and scalability of the SCER framework. These experiments provide detailed insights into SCER’s adaptation mechanisms across various scenarios and assess its practical feasibility for real-world deployment.
@@
While our method requires computational investment for correlation metric and mean embedding computations, it maintains computational costs comparable to established robust training methods. As shown in Table C.1, our approach requires 0.0988 seconds per iteration, which is marginally higher than GroupDRO at 0.0915 seconds but significantly faster than the two-stage DFR method at 0.1025 seconds.
Importantly, DFR’s computational overhead stems from its inherent two-stage paradigm (kirichenko2022last): first training with ERM to learn representations, then retraining only the final layer with group-balanced data. This approach suffers from critical limitations, including the requirement for high-quality, balanced validation data and complete failure when ERM cannot capture core features, such as when minority groups are entirely absent from training data. In contrast, our single-stage approach achieves substantially better worst-group accuracy than these alternatives while avoiding DFR’s structural constraints. Most importantly, our method requires no balanced validation data or multistage training, delivering superior robustness with greater practical simplicity.

### v2

@@ B.1 Experimental Setting
This section provides comprehensive details of our experimental configuration for evaluating the SCER framework. We describe the datasets used for evaluation, the model architectures employed across different domains, the baseline methods used for comparison, the hyperparameter settings, and the evaluation metrics. These detailed specifications ensure reproducibility and provide a complete context for the results presented in the main paper.
@@ B.1.1 Datasets
We conduct experiments on six benchmark datasets. Four are image datasets, including Waterbirds, CelebA, MetaShift, and ColorMNIST, and two are text datasets, including CivilComments and MultiNLI. These datasets are established benchmarks for evaluating model robustness against spurious correlations. Dataset characteristics and statistics are summarized in Table 8.
Table 8: Statistics of datasets. For ColorMNIST, $\rho$ denotes the spurious correlation level between label $y$ and domain attribute (color) in the training set. A higher $\rho$ indicates stronger spurious correlation, making color a more predictive but misleading cue for classification. ”One Subpopulation Omitted” represents an extreme case in which one group combination is completely absent from the training data. Groups are defined as combinations of class labels and attribute values, and Max/Min group columns show the sample sizes of the largest and smallest groups, respectively. The disparity between these values indicates the degree of group imbalance, which tests model robustness under uneven data distribution.
@@
Waterbirds (Sagawa et al., ) is a binary classification dataset constructed by superimposing bird images from the CUB dataset (Wah et al., 2011) onto backgrounds from the Places dataset (Zhou et al., 2017). The task involves classifying images as landbirds or waterbirds, where background type creates spurious correlations between land backgrounds and landbirds, and water backgrounds and waterbirds. We use the standard data splits from Idrissi et al. (2022).
@@
CelebA (Liu et al., 2015) contains approximately 200,000 celebrity images. We formulate the problem as a binary classification task, classifying hair as either blond or non-blond, with gender serving as the spurious attribute. We follow the standard splits from Idrissi et al. (2022).
@@
From the MetaShift benchmark (Liang and Zou, 2022), we use the Cats vs. Dogs binary classification task, where background type serves as the spurious attribute through correlations between indoor backgrounds with cats and outdoor backgrounds with dogs. We adopt the unmixed version provided by the dataset authors.
@@
ColorMNIST (Arjovsky et al., 2019) is a synthetic variant of MNIST where each digit is assigned a color. The classification task groups digits into two categories: zero and four versus five and nine, with color serving as the spurious attribute through red and green assignments. Training data exhibits strong color-class correlations, validation data is balanced at 50:50, and test data reverses the correlation pattern to 10:90. Following Arjovsky et al. (2019), we inject 25% label noise and evaluate under varying spurious correlation strengths of 80:20, 90:10, and 95:5 ratios to assess robustness.
To test the limits of group imbalance, we conducted additional experiments on ColorMNIST with severe distributional skew, in which one subpopulation is completely omitted from training. Specifically, we use 15,249 samples for Label 0 with Color 0, 3,000 samples for Label 1 with Color 0, and 11,751 samples for Label 1 with Color 1, while Label 0 with Color 1 contains zero samples. This extreme setting evaluates model performance when specific subpopulations are entirely missing during training, representing a challenging scenario for robust classification methods.
To test the generalizability of SCER, we conducted additional experiments on ColorMNIST using the environment inference methodology from Creager et al. (2021). Following the prior protocol with two environments in train, we partition the dataset into 50,000 training samples and 10,000 test samples. The training set is further divided into two environments, each containing 25,000 samples that represent different spurious correlation patterns. We maintain consistency with our primary experimental setup by using identical model architectures, training procedures, and evaluation metrics across all experiments.
@@
CivilComments (Borkan et al., 2019) provides a binary text classification benchmark designed to identify toxic language in online comments. Mentions of demographic groups serve as spurious attributes. We use the standard WILDS splits.
CivilComments (4 groups)
As an additional experiment on the CivilComments (Borkan et al., 2019) dataset, we use the coarse version for both training and evaluation to enable a fair comparison with GroupDRO-ES (Izmailov et al., 2022), which is known as the most competitive GroupDRO baseline and represents the state-of-the-art GroupDRO performance. In this setting, the spurious attribute is defined as 1 if the comment mentions at least one of the following categories, including male, female, LGBT, black, white, Christian, Muslim, or other religion. Otherwise, the spurious label is 0. The presence of these eight categories is spuriously correlated with the comment being classified as toxic. This configuration forms four groups, with two classes for toxic and non-toxic comments, and two attributes representing the presence or absence of the spurious attribute.
@@
The MultiNLI dataset (Williams et al., 2018) addresses natural language inference by requiring models to classify the relationship between a premise and a hypothesis as entailment, contradiction, or neutrality. The presence of negation words serves as a confounding factor because they exhibit a strong correlation with contradiction labels. We adopt the data splits established in Idrissi et al. (2022).
@@
Table 10: Summary of attribute availability settings across baseline methods.
Train Attr. Valid/Held-out Attr. Method
Known Known GroupDRO, LISA, CVaRDRO, ReSample, PDE, ElRep(with GroupDRO)
Unknown Known (Val) JTT, LfF, CRT, SqrtReWeight
Unknown Known (Held-out) DFR
Unknown Unknown ReWeight, ERM, Mixup, MMD, Focal Loss, CB Loss, LDAM, Balanced Softmax
B.1.2 Baselines
For fair comparison, we followed the training protocol of Gulrajani and Lopez-Paz , performing random search over each method’s joint hyperparameter space. Hyperparameter configurations follow prior work (Yang et al., 2023b) for all baseline methods. We selected the hyperparameters that achieved the highest Worst Accuracy for each algorithm, then conducted three independent runs with different random seeds (0, 1, 2) to account for variance and ensure robustness of reported results. Final average performance and standard deviation are reported across these runs. Complete hyperparameter details are provided in Table 9, which summarizes the hyperparameters used for each baseline. In addition to standard optimization parameters such as batch size, learning rate, and dropout, we include key weighting parameters $\eta$, $\lambda_{spur}$, and $\lambda_{core}$ to suppress spurious features and enhance core features. These values are chosen to encourage the model to learn generalizable representations that generalize across distribution shifts.
We provide a clear summary in Table 10  indicating which baselines utilize group annotations at each stage: training, held-out selection, and validation. Following the evaluation protocol established in prior work (Yang et al., 2023b), we adopt the Known Attributes setting, in which group annotations are available for both the training and validation sets. All baselines are evaluated under this unified setting to ensure consistency and fair comparison. While some methods are designed to operate without group annotations during training or validation, all methods ultimately rely on group information for final model selection. Specifically, we select the best checkpoint for each method based on worst-group accuracy, ensuring alignment in model selection across all approaches. Thus, even when a method does not internally use group information during optimization, group annotations are consistently used for model selection to ensure fair and comparable evaluation.
To comprehensively evaluate our method’s robustness, we compare its performance against a diverse set of baselines spanning standard ERM, subpopulation robustness methods, domain-invariant approaches, data augmentation techniques, and class imbalance methods. The primary baseline results appear in Tables 125.1, and 5.1 in the main text, with complete results for all baselines provided in Tables 1112C.1 and C.1. Our baseline selection follows evaluation criteria established in recent studies (Deng et al., 2023; Wen et al., 2025), with particular attention to whether each method utilizes group annotations during training.
To ensure fair comparison, the main text reports only methods that rely on group annotations throughout the entire training and validation pipeline, while methods not satisfying this criterion are presented separately in the appendix to reflect differences in evaluation settings.
@@ B.1.2 Baselines
GroupDRO (Sagawa et al., ) is a representative approach to distributional robustness that improves worst-group performance by assigning greater importance to groups with higher losses. CVaRDRO (Duchi and Namkoong, 2021) extends this approach to the individual sample level, enabling more fine-grained adjustment by assigning importance weights to individual samples. LfF (Nam et al., 2020) employs a co-training approach that simultaneously trains a biased model and a debiased model, guiding the model to learn by distinguishing between biased and unbiased features. JTT (Liu et al., 2021) is a two-stage strategy that identifies complex samples using information obtained from initial standard empirical risk minimization (ERM) training, then improves model robustness by increasing the weights of these samples in the second training stage. LISA (Yao et al., 2022) utilizes data augmentation techniques that mix data within and across attributes, encouraging the model to learn invariant features of attributes. DFR (Kirichenko et al., ) is an efficient method that improves robustness by retraining only the last layer using a balanced validation set, rather than the entire model. PDE (Deng et al., 2023) employs a strategy of progressively expanding training data, guiding the model to first focus on core features and subsequently learn spurious features. Finally, ElRep (Wen et al., 2025) introduces both Nuclear-norm and Frobenius-norm regularization to representation vectors, enabling robust group generalization representation learning against spurious correlations by simultaneously emphasizing important features and maintaining diversity.
@@
IRM (Arjovsky et al., 2019) is a method for finding invariant predictors that work commonly across multiple domains. Specifically, it trains feature extractors so that a fixed predictor becomes optimal in each domain, enabling the model to capture important characteristics effectively regardless of domain changes. This approach ensures robust performance even when domains change. MMD (Li et al., 2018) is a method that numerically measures the difference in feature distributions across domains and trains to reduce this difference, thereby mitigating distribution mismatches. In other words, it is an approach that aims to improve generalization performance by making feature representations across domains as similar as possible.
@@
Mixup (Zhang et al., 2018) is a technique that generates synthetic training data by linearly interpolating two randomly sampled pairs of inputs and their labels. This smooths decision boundaries and serves as regularization to improve generalization performance and prevent overfitting. It is a simple yet effective data augmentation method that artificially expands the diversity of the training data by learning intermediate points between samples in the data space.
@@
ReSample (Japkowicz, 2000) mitigates class imbalance in training data through resampling techniques. ReWeight (Japkowicz, 2000) assigns weights to the loss function proportional to class frequency. SQRTWeight is another loss adjustment method that adjusts loss weights using the inverse square root of class frequencies, thereby increasing the influence of minority classes in imbalanced data. Focal Loss (Lin et al., 2017) focuses on difficult samples by reducing the contribution of easy-to-learn samples. CB Loss (Cui et al., 2019) calculates sample weights based on effective class size. LDAM (Cao et al., 2019) assigns distinct margins to each class to balance the decision boundaries. Balanced Softmax (Ren et al., 2020) modifies the softmax function itself to accommodate class imbalance.
Table 11: Complete performance comparison on multiple datasets, each exhibiting different levels of spurious correlations. denotes the performance reported from Yang et al. (2023b) except for ColorMNIST. denotes the performance reported from Deng et al. (2023) for Waterbird and CelebA. †† denotes the performance reported from Wen et al. (2025) for Waterbird and CelebA. For the remaining datasets, the detailed experimental settings follow Yang et al. (2023b), as described in Appendix B.1.2. SCER achieves the highest worst-group accuracy in Waterbirds, CelebA, MetaShift, and ColorMNIST, demonstrating its robustness in subpopulation shift scenarios.
Algorithm Waterbirds CelebA MetaShift ColorMNIST ($\rho=80\%$)
Avg Acc Worst Acc Avg Acc Worst Acc Avg Acc Worst Acc Avg Acc Worst Acc
[…]
@@ C.1 Quantitative Analysis
Here, we present the complete results for quantitative analyses omitted from the main text.
Table 12: Complete performance comparison on ColorMNIST under increasing spurious correlation levels. SCER achieves the highest worst-group accuracy across different spurious correlation levels, demonstrating robustness to spurious feature reliance in ColorMNIST.
ColorMNIST
$\rho=80\%$ $\rho=90\%$ $\rho=95\%$ $\rho=99\%$
Algorithm Avg Acc Worst Acc Avg Acc Worst Acc Avg Acc Worst Acc Avg Acc Worst Acc
ERM 38.2 $\pm$2.4 30.9 $\pm$2.7 31.9 $\pm$7.4 7.8 $\pm$3.4 41.1 $\pm$13.0 10.1 $\pm$4.5 50.7 $\pm$8.2 8.5 $\pm$5.0
GroupDRO 73.5 $\pm$0.3 73.1 $\pm$0.1 73.5 $\pm$0.0 72.7 $\pm$0.3 72.3 $\pm$0.1 70.7 $\pm$0.2 50.3 $\pm$3.9 38.7 $\pm$0.8
LISA 73.8 $\pm$0.1 73.2 $\pm$0.1 73.3 $\pm$0.1 72.9 $\pm$0.2 72.6 $\pm$0.4 71.4 $\pm$0.6 53.1 $\pm$7.2 10.2 $\pm$8.1
ReSample 73.7 $\pm$0.1 72.2 $\pm$0.3 72.6 $\pm$0.5 70.8 $\pm$1.6 71.8 $\pm$0.1 70.7 $\pm$0.5 57.1 $\pm$3.0 45.4 $\pm$10.0
PDE 76.6 $\pm$0.8 72.9 $\pm$0.3 72.6 $\pm$0.2 70.2 $\pm$0.7 73.4 $\pm$0.4 70.0 $\pm$0.3 56.3 $\pm$4.3 47.5 $\pm$8.9
SCER 74.1 $\pm$0.1 73.6 $\pm$0.2 73.6 $\pm$0.1 73.0 $\pm$0.1 73.5 $\pm$0.3 72.8 $\pm$0.3 61.0 $\pm$2.0 56.0 $\pm$2.2
Table 13: Complete performance on ColorMNIST with one group absent. SCER achieves the highest worst-group accuracy.
ColorMNIST (One Subpopulation Omitted)
Algorithm Avg Acc Worst Acc
ERM 52.9 $\pm$8.4 9.7 $\pm$5.4
GroupDRO 53.6 $\pm$2.5 44.1 $\pm$6.2
LISA 49.5 $\pm$2.4 13.5 $\pm$11.0
ReSample 49.1 $\pm$0.8 16.3 $\pm$13.0
ReWeight 52.1 $\pm$4.7 10.0 $\pm$8.2
CBLoss 53.0 $\pm$3.2 13.3 $\pm$10.0
[…]
SCER 65.3 $\pm$1.0 59.6 $\pm$1.0
Image Dataset
Table 11 reports the complete results for all baselines on image datasets. Across this comprehensive set of methods, SCER consistently achieves the highest worst-group accuracy, demonstrating its superior robustness compared to all competing approaches.
Image Dataset with Stronger Spurious Correlations
Table 12 reports the complete results on ColorMNIST under varying spurious correlation levels, including all methods achieving worst-group accuracy above 70% in Table 11. SCER achieves the highest worst-group accuracy across most settings, confirming its robustness under subpopulation shift.
In addition to the standard evaluation settings in Table 12, we examine a more challenging scenario in which a minority group is completely absent during training. Table C.1 reports the complete results under this extreme setting. As shown, SCER achieves the highest worst-group accuracy, significantly outperforming all competing methods. Notably, class imbalance methods such as ReWeight (Japkowicz, 2000) and CB Loss (Cui et al., 2019) exhibit substantial performance degradation when a group is entirely missing, similar to other approaches. In contrast, SCER maintains robust performance even under this extreme distributional shift, demonstrating its effectiveness in handling severe group imbalance.
Table 14: Performance comparison without explicit bias labels in noisy environments.
ColorMNIST (Two train envs)
Method Avg Acc
IRM 54.8 $\pm$8.3
EIIL (IRM) 53.7 $\pm$2.1
EIIL + DRO 54.6 $\pm$5.9
EIIL + SCER 65.1 $\pm$2.8
Integration with Environment Inference Methods.
Following Table 5.1, we additionally experimented to verify whether SCER maintains robustness when the environment partitioning accuracy is imperfect, at approximately 70%. All other settings remain identical to those in Table 5.1, and the results are presented in Table C.1. Even under reduced environment partitioning accuracy, EIIL + SCER achieves an average accuracy of 65.1%, outperforming both EIIL + DRO at 54.6% and EIIL with IRM at 53.7%. This suggests that SCER operates effectively even under imperfect environment partitioning conditions.
Text Dataset
Table C.1 presents the complete performance comparison on text datasets, including CivilComments and MultiNLI. SCER consistently achieves superior worst-group accuracy across these benchmarks, demonstrating its effectiveness beyond image domains. On CivilComments, SCER attains a worst-group accuracy of 74.0% . Similarly, on MultiNLI, SCER achieves the best worst-group accuracy of 76.8%.
Tables 1112C.1, and C.1 demonstrate that SCER consistently achieves superior performance across diverse benchmark datasets and in comparison with various state-of-the-art algorithms. Notably, SCER achieves robust performance not only on vision tasks but also on natural language understanding tasks, confirming that our mechanism of separating spurious and core features at the embedding level generalizes effectively across domains.
Table 15: Complete performance comparison on text datasets. from Deng et al. (2023) for Civilcomments. †† from Wen et al. (2025) for Civilcomments. SCER achieves the best worst-group accuracy across multi-class and domain settings.
Algorithm CivilComments MultiNLI
Avg Acc Worst Acc Avg Acc Worst Acc
ERM 85.4 $\pm$0.2 63.7 $\pm$1.1 80.9 $\pm$0.1 66.8 $\pm$0.5
Mixup 84.9 $\pm$0.3 66.1 $\pm$1.3 81.4 $\pm$0.3 68.5 $\pm$0.6
GroupDRO 81.8 $\pm$0.6 70.6 $\pm$1.2 81.1 $\pm$0.3 76.0 $\pm$0.7
IRM 85.5 $\pm$0.0 63.2 $\pm$0.8 77.8 $\pm$0.6 63.6 $\pm$1.3
CVaRDRO 83.5 $\pm$0.3 68.7 $\pm$1.3 75.1 $\pm$0.1 63.0 $\pm$1.5
JTT 83.3 $\pm$0.1 64.3 $\pm$1.5 80.9 $\pm$0.5 69.1 $\pm$0.1
LfF 65.5 $\pm$5.6 51.0 $\pm$6.1 71.7 $\pm$1.1 63.6 $\pm$2.9
LISA 82.7 $\pm$0.1 73.7 $\pm$0.3 80.3 $\pm$0.4 73.3 $\pm$1.0
MMD 84.6 $\pm$0.2 54.5 $\pm$1.4 78.8 $\pm$0.1 69.1 $\pm$1.5
ReSample 82.2 $\pm$0.0 73.3 $\pm$0.5 77.2 $\pm$0.2 72.3 $\pm$0.8
ReWeight 82.5 $\pm$0.0 72.5 $\pm$0.0 81.0 $\pm$0.2 68.8 $\pm$0.4
SqrtReWeight 83.3 $\pm$0.5 71.7 $\pm$0.4 80.7 $\pm$0.3 69.5 $\pm$0.7
CBLoss 82.9 $\pm$0.1 73.3 $\pm$0.2 80.6 $\pm$0.1 72.2 $\pm$0.3
Focal 85.5 $\pm$0.2 62.0 $\pm$1.0 80.7 $\pm$0.2 69.4 $\pm$0.7
LDAM 81.9 $\pm$2.2 37.4 $\pm$8.1 80.7 $\pm$0.3 69.6 $\pm$1.6
BSoftmax 83.8 $\pm$0.0 71.2 $\pm$0.4 80.9 $\pm$0.1 66.9 $\pm$0.4
DFR 83.3 $\pm$0.0 69.6 $\pm$0.2 81.7 $\pm$0.0 68.5 $\pm$0.2
PDE 86.3 $\pm$1.7 71.5 $\pm$0.5 69.1 $\pm$0.3 65.8 $\pm$0.6
ElRep†† 79.0 $\pm$0.7 70.5 $\pm$0.5 69.0 $\pm$3.8 66.8 $\pm$5.2
SCER 81.7 $\pm$0.4 74.0 $\pm$1.0 80.4 $\pm$0.1 76.8 $\pm$0.5
Table 16: Performance comparison on Waterbirds, CelebA, MultiNLI, and CivilComments (4 groups) with GroupDro-ES. § from Izmailov et al. (2022). Experimental and dataset settings also follow Izmailov et al. (2022). SCER achieves the best worst-group accuracy across all datasets.
Algorithm Waterbirds CelebA MultiNLI CivilComments (4 groups)
Worst Acc Worst Acc Worst Acc Worst Acc
GroupDRO-ES§ $90.7\pm 0.6$ $90.6\pm 1.6$ $73.5$ $80.4$
SCER $\mathbf{91.7\pm 0.1}$ $\mathbf{91.9\pm 0.3}$ $\mathbf{77.9}$ $\mathbf{82.9}$
Deep Analysis with GroupDRO
SCER proposes a novel objective function that decomposes worst-group error into spurious and core components, then directly regularizes the embedding space to suppress spurious correlations while enhancing core features. To validate the effectiveness of this approach, we conducted an in-depth comparison with GroupDRO. Since early stopping is well known to be essential for achieving optimal GroupDRO performance (Izmailov et al., 2022), we compared SCER with GroupDRO-ES (Izmailov et al., 2022), which serves as the most competitive GroupDRO baseline. Note that for a fair and direct comparison with GroupDRO-ES, we adopted the experimental setup from Izmailov et al. (2022) and used their reported results for GroupDRO-ES, which differ from the configuration in our main experiments following Yang et al. (2023b).
As shown in Table 16, SCER achieves consistent performance improvements over GroupDRO-ES across all datasets. Specifically, SCER demonstrates gains of 1.0% on Waterbirds and 1.3% on CelebA, with a particularly notable 4.4% improvement on MultiNLI. This represents meaningful progress, given that most group-robust methods struggle to exceed worst-group accuracy in the low 70% range on this dataset. For CivilComments, while Tables 5.1 and C.1 follow the 16 group setting with two classes and eight attributes as in prior work (Yang et al., 2023b), we also conducted experiments using the 4 group setting with two classes and two attributes in Table 16 to allow a fair comparison with (Izmailov et al., 2022), achieving a 2.5% improvement. Detailed information about the dataset configurations is provided in Table 8.
Beyond quantitative gains, we provide qualitative evidence that SCER fundamentally restructures the embedding space. Figure 3 shows through t-SNE analysis that while ERM and GroupDRO separate clusters by background, SCER achieves label-based clustering independent of spurious features, confirming domain-invariant representations. Figure 5 further demonstrates via Grad-CAM that SCER attends to target objects rather than backgrounds, showing that our embedding regularization yields genuine improvements in representation learning.
@@
We also conduct a comprehensive quantitative analysis to evaluate the computational complexity and scalability of the SCER framework. These experiments provide detailed insights into SCER’s adaptation mechanisms across various scenarios and assess its practical feasibility for real-world deployment.
@@
While our method requires computational investment for correlation metric and mean embedding computations, it maintains computational costs comparable to established robust training methods. As shown in Table C.1, our approach requires 0.0988 seconds per iteration, which is marginally higher than GroupDRO at 0.0915 seconds but faster than the two-stage DFR method at 0.1025 seconds.
Importantly, DFR’s computational overhead stems from its inherent two-stage paradigm (Kirichenko et al., ): first training with ERM to learn representations, then retraining only the final layer with group-balanced data. This approach suffers from critical limitations, including the requirement for high-quality, balanced validation data and complete failure when ERM cannot capture core features, such as when minority groups are entirely absent from training data. In contrast, our single-stage approach achieves substantially better worst-group accuracy than these alternatives while avoiding DFR’s structural constraints. Most importantly, our method requires no balanced validation data or multistage training, delivering superior robustness with greater practical simplicity.